=== 2019/2/1(Fri.) ===

=== Weekend ===
 * Alignment work (Sat and Sun): Nakano, Enomoto, Guiguo, Wang
 * Cal: ETMY Inoue, Ohgaki
 * Miyoki(Sat)

=== Schedule ===
  * Current status as of 1/22: IR and Green reached X end PD.
  * --(1/22: Surface inspection for PR mirrors.)-- Cleaning will be done after 1/31.
  * --(1/23: Surface inspection for SR mirrors.)--
  * --(1/24-25: Cleaning HR of SR3, SRM.)--

  * --(1/23: BS is aligned to Y-end PD by IR.)--
  * --(1/27: ITMY is aligned to REFL by IR.)--
  * --(1/27: ITMX is aligned to REFL by IR.)--
  * --(1/27: Beam position of IR at SR3 is confirmed how far from the center.)--

  * --(1/28-29: SR3 is aligned to SR2 center by IR.)--
  * --(1/28: Alignment of SR2 -> SRM by IR.)--
  * --(1/29: Green beam is injected from SR2 AR surface to Y end PD.)--

  * --(1/31: Removing bellows at PR and SR area.)--
  * 1/31-2/1: Cleaning HR of PR3 and HR, AR of SR2

  * 2/4-8: SRM -> OMMT by IR -> OMMT -> OSTM by IR, OSTM -> OMC by IR
  * 2/4-8: Alignment of ETMY -> POS by Green

  * 2/12-15: Closing door for central area, pumping down
   * needs drift monitor for VIS

==== Note ====

 * Please report everything you did to klog, and gokagra ML if urgent things.

==== Report of work ====
 * VAC: Bellows were removed at PR3 and SR2.
 * MIR: FC was put on HR of PR3 and AR of SR2.
  * FC of HR of SR will be put on today.
  * All FC will be removed today.
 * IOO: MZ was locked. we will do some measurement. Clipping was fixed.
 * Type A: ITMX is being tuned continuously.
 * PEM: ETMX fan is stopped now.
 * ETMY: we found some dust on the surface. We will talk with MIR group.


==== Work plan ====
 * High power test in the weekend 
  * Laser comes to IFI with high power.
  * Laser comes to IFI today. Beam will be stopped at IMMT.
 * Type B: View port on SRM will be installed. FR of SR and some miner issues will be fixed.
  * Height change due to FC will be fixed today.